### v4

@@ Abstract
Keywords:
Log Parsing Large Language Models Information Entropy.
@@ 1 Introduction
Logs serve as a critical information source for system monitoring, offering key insights into system behavior, a fact well-documented in existing literature [12]. Their advantage over other types of data lies in their rich informational content and relative ease of interpretation. Through log analysis, several important downstream tasks can be effectively addressed, which include anomaly detection [5, 23], fault diagnosis [13, 39], and root cause analysis [31]. Log parsing, a crucial initial step in log analysis [10], separates log messages into two parts: 1) Log Templates. The constant, unchanging parts in logging statements, and 2) Log Variables. The dynamic, changeable details in different executions. In Figure 1, the logging statement logger.info(’Wait {wait_time} seconds for {process} to be killed.’) can yield various messages like Wait 26 seconds for Thread-20 to be killed. Wait <><*>< ∗ > seconds for <><*>< ∗ > to be killed is the log template, and the changing data like 26 and Thread-20 are the log variables.
@@
In the field of system maintenance, where source code accessibility is often restricted, different parsers [11, 7, 18, 38, 15] have been developed to differentiate between templates and variables within log messages autonomously. Existing log parsers can be categorized into two groups: syntax-based and semantic-based. Syntax-based log parsers [11, 4, 21, 32, 33] utilize specific features or heuristics (e.g., word frequency) to extract the constant parts of log messages as templates. Conversely, semantic-based parsers [20, 18] leverages advanced deep learning frameworks to assimilate semantics and system-specific patterns from annotated log data, thereby facilitating the parsing of new logs. Besides, recent works [15] leverage large language models (LLMs) [25] for log parsing, which also utilizes the powerful semantic understanding of LLMs.
@@
However, syntax-based methodologies are heavily dependent on meticulously crafted rules. Their performance significantly diminishes with the exponential increase in the volume and intricacy of log data. Furthermore, these techniques often overlook semantic variances among logs. For instance, logs such as success to open file/var/log/system and success to close file/var/log/system display syntactic resemblance, potentially leading to their aggregation by data-driven approaches, which could adversely impact the efficacy of anomaly detection tasks. Semantic-based methods, reliant on pre-labeled logs for training, falter in understanding semantics when applied to unfamiliar domains. LLM-based parsers [19, 18, 15] have lower availability due to high inference time (such as GPUs) and network latency [16]. Besides, LLMs generate unstable results because of the hallucination problem.
@@
To tackle these challenges, we propose Lemur, a cutting-edge Log parsing framework with Entropy sampling and Chain-of-Thought Merging (Lemur). Lemur brings together the strengths of the syntax-based and semantic-based methods, which consist of three key components: Information Entropy Clustering, Template Generation, and Chain-of-Thought Merging. Specifically, inspired by information theory [8], we recognize that different logs encapsulate varying quantities of information, while logs of a similar nature contain comparable amounts. Consequently, we have developed a novel sampling method based on information entropy principles that efficiently clusters characteristic logs by dividing large-scale data into multiple clusters and utilizing efficient sampling and clustering algorithms within those clusters, thereby ensuring that LEMUR remains robust and high-performance in large-scale log scenarios. Then in template generation, we determine the variables and the template in the log based on the information entropy of the token at the same location. In the end, motivated by Chain-of-Thought [35], we design a three-hop Chain-of-Thought (infer structure, infer semantics, and infer solution) for merging templates.
@@
We have conducted a comprehensive evaluation on public large-scale log datasets of LogHub [14] with seven state-of-the-art log parsing approaches. The results show that Lemur achieves the highest performance than other baselines for the F1 score of grouping and template accuracy. Generally, the main contributions of this work are listed as follows:
@@ 2.1 Information Entropy Clustering
To identify potential log templates from buckets of logs with identical token lengths, we propose a clustering method inspired by information theory [8], specifically employing Shannon entropy. Unlike conventional clustering which uses random seeds and iterative refinement, our approach selects logs based on their information content, which is ideal for log analysis due to the variability in log messages.
@@ 2.3 Three-S-hop Chain-of-Thought Merging
Figure 3: Three-S-hop Chain-of-Thought Merging Method
@@
Dataset Template Avg. Len. Various Lengths Proportion Messages
HDFS 14 7.44 0.25% 2000
Hadoop 114 8.19 0.00% 2000
Spark 36 8.76 0.00% 2000
Zookeeper 50 6.3 0.55% 2000
BGL 120 6.32 0.00% 2000
HPC 46 3.48 0.40% 2000
Thunderbird 149 8.51 2.25% 2000
Windows 50 7.93 0.00% 2000
Linux 118 8.3 0.00% 2000
Andriod 166 5.4 3.85% 2000
HealthApp 75 2.8 0.00% 2000
Apache 6 6.28 0.00% 2000
Proxifier 8 9.35 47.35% 2000
OpenSSH 27 8.56 0.00% 2000
OpenStack 43 9.01 0.00% 2000
Mac 341 9.17 0.35% 2000
@@
Dataset $D$ $k$ $J_{T}$ $\theta$
HDFS : 2 0.7 2.0
Hadoop =, :, (, ), _ 8 0.7 1.7
Spark : 6 0.6 2.1
Zookeeper =, :, , 8 0.9 2.2
BGL =, ., ., (, ) 9 0.6 5.5
HPC =, -, : 9 0.6 1.2
Thunderbird :, = 11 0.4 4.1
Windows =, :, [, ] 8 0.6 1.1
Linux =, : 25 0.33 0.09
Andriod (, ) 9 0.7 3.5
HealthApp =, : 12 0.7 0
Apache 12 0.7 0
Proxifier 12 0.7 0.1
OpenSSH = 4 0.5 0.2
OpenStack 20 0.7 2.3
Mac 12 0.7 4.7
Table 2: Hyperparameters setting of Loghub datasets. $D$ is the tokens for word-splitting, $k$ denotes the number of the $S$ in each bucket, $J_{T}$ is the Jaccard similarity threshold for within-bucket merging, and $\theta$ is the entropy threshold for token identification.
@@ 3.1 Implement Details
Experiments are conducted on the most widely-used benchmark datasets published in LogHub[14]. More details are available in Table 1.
@@ 3.2 Baselines and Metrics
As for baselines, we choose Drain[11], Spell[4], IPLOM[21], ULP[28], Brain[38], LogPPT[18] and LLMParser[15] as our baselines.
@@ 3.3 Main Results
Figure 4: Boxplot of group accuracy on 16 benchmark datasets.
Figure 5: Boxplot of F1 score of group accuracy on 16 benchmark datasets.
@@
3.4 Effect of Chain-of-Thought Merging
Dataset Lemur w/o CoT Lemur
FGA GA FGA GA
HDFS 0.999 0.998 1 1
Hadoop 0.999 0.977 1 1
Spark 1 1 1 1
Zookeeper 0.999 0.995 1 1
BGL 0.999 0.989 1 1
HPC 0.999 0.996 1 1
Thunderbird 0.999 0.959 0.999 0.982
Windows 1 1 1 1
Linux 0.999 0.986 0.999 0.988
Andriod 0.992 0.956 0.999 0.995
HealthApp 1 1 1 1
Apache 1 1 1 1
Proxifier 1 1 1 1
OpenSSH 1 1 1 1
OpenStack 1 1 1 1
Mac 0.982 0.956 0.984 0.977
Average 0.998 0.988 0.999 0.996
Table 5: FGA and GA on LogHub bewteen Lemur with and without CoT.
@@ 3.4 Effect of Chain-of-Thought Merging
KMeans Clustering [30] 0.685 0.698
DBSCAN Clustering [17] 0.731 0.744
@@
In Figure 5 and Figure 4, Lemur shows the robust performance in GA and FGA. The consistency and robustness of Lemur are evident in its high FGA and GA across diverse datasets such as HDFS, Hadoop, Spark, etc. This consistency emphasizes its adaptability and robustness to various log parsing requirements. In conclusion, as an unsupervised LLM-related model, Lemur significantly stands out in the log parsing domain.
@@ 4.2 Additional Ablation Studies on Lemur
To further validate the robustness of Lemur, we conducted ablation studies focusing on the impact of different clustering algorithms and the effectiveness of the bucket grouping strategy. Specifically, we experimented with replacing our clustering mechanism with KMeans [30] and DBSCAN [17] and evaluated the performance when disabling the bucket grouping based on the length assumption. These modifications aim to probe the sensitivity of Lemur to these components.
@@ 5.1 Log Parser
In the evolving field of automatic log analysis, crucial for distributed systems and cloud computing, significant progress has been made in log parsing techniques, categorized into frequent pattern mining, clustering, and heuristics rules. Frequent pattern mining is exemplified by SLCT [32] which groups logs based on token frequency, and LogCluster [33] which removes positional constraints in log grouping. Clustering approaches include LogMine [9] with its multi-layered clustering system, LKE[6] using edit distance and position weighing, SHISO [22] improving efficiency through hierarchical clustering, LenMa [29] employing token length vectors, and LPV [37] which uses semantic vectors from word2vec. In the heuristics rules category, IPLOM [21] partitions logs by length and token position, Spell [4] approaches parsing as the longest common sequential problem, Drain [11] groups logs by length and prefixes for template updates, and Prefix-Graph [3] merges prefix trees into graphs for template generation. Recent advancements have introduced deep learning-based algorithms like Nulog [24] Uniparser [20], and LogAP [27] utilizing comparative learning and machine translation for parsing. However, these deep learning methods face challenges in efficiency and high operational costs due to GPU requirements.
@@ 5.2 Large Language Model
Language modeling using self-supervised learning and large-scale data, significantly enhances various natural language processing tasks. Specifically, pre-training a Transformer decoder[25, 34] aids in unconditional text generation. Performance improvements in diverse tasks have been linked to the enlargement of Pre-training Language Models (PLMs) by increasing model or data size. This has led to the creation of increasingly larger PLMs, such as GPT-3 with 175 billion parameters and PaLM with 540 billion [2], guided by the scaling laws of large language models [1]. Despite their similar architectures and pre-training tasks, larger PLMs, such as GPT-4 [25], exhibit unique behaviors and emergent abilities, excelling in complex tasks. A prime example is ChatGPT, adapting GPT-series LLMs for engaging dialogues, and showcasing advanced conversational skills. Fine-tuning LLMs on various datasets [36] yields promising results, using human or LLM-created prompts for instruction tuning and refining generations. Chain-of-thought prompting [36], where models explain their reasoning for complex problems, and RLHF [26], a strategy significantly enhance their performance.

### v5

@@ 1 Introduction
Logs serve as a critical information for system monitoring, offering key insights into system behavior, a fact well-documented in existing literature (He et al., 2021). Their advantage over other types of data lies in their rich informational content and relative ease of interpretation. Through log analysis, several important downstream tasks can be effectively addressed, which include anomaly detection (Du et al., 2017; Nandi et al., 2016), fault diagnosis (He et al., 2018; Zou et al., 2016), and root cause analysis (Tak et al., 2016). Log parsing, a crucial initial step in log analysis (He et al., 2016), separates log messages into two parts: 1) Log Templates. The constant, unchanging parts in logging statements, and 2) Log Variables. The dynamic, changeable details in different executions. In Figure 1, the logging statement logger.info(’Wait {wait_time} seconds for {process} to be killed.’) can yield various messages like Wait 26 seconds for Thread-20 to be killed. Wait <><*>< ∗ > seconds for <><*>< ∗ > to be killed is log template, and the changing data like 26 and Thread-20 are log variables.
@@
In the field of system maintenance, where source code accessibility is often restricted, different parsers (He et al., 2017; Fu et al., 2022; Le & Zhang, 2023; Yu et al., 2023; Jiang et al., 2023) have been developed to differentiate between templates and variables within log messages autonomously. Existing log parsers can be categorized into two groups: syntax-based and semantic-based. Syntax-based log parsers (He et al., 2017; Du & Li, 2016; Makanju et al., 2009; Vaarandi, 2003; Vaarandi & Pihelgas, 2015) utilize specific features or heuristics (e.g., word frequency) to extract the constant parts of log messages as templates. Conversely, semantic-based parsers Liu et al. (2022); Le & Zhang (2023) leverages advanced deep learning frameworks to assimilate semantics and system-specific patterns from annotated log data, thereby facilitating the parsing of new logs. Besides, recent works Jiang et al. (2023) leverage large language models (LLMs) (OpenAI, 2023) for log parsing, which also utilizes the powerful semantic understanding of LLMs.
@@
However, syntax-based methodologies are heavily dependent on meticulously crafted rules. Their performance significantly diminishes with the exponential increase in the volume and intricacy of log data. Furthermore, these techniques often overlook semantic variances among logs. For instance, logs such as success to open file/var/log/system and success to close file/var/log/system display syntactic resemblance, potentially leading to their aggregation by data-driven approaches, which could adversely impact the efficacy of anomaly detection tasks. Semantic-based methods, reliant on pre-labeled logs for training, falter in understanding semantics when applied to unfamiliar domains. LLM-based parsers (Liu et al., 2023; Le & Zhang, 2023; Jiang et al., 2023) have lower availability due to high inference time (such as GPUs) and network latency (Jiao et al., 2023). Besides, LLMs generate unstable results because of the hallucination problem.
@@
To tackle these challenges, we propose Lemur, a cutting-edge Log parsing framework with Entropy sampling and Chain-of-Thought Merging (Lemur). Lemur brings together the strengths of the syntax-based and semantic-based methods, which consist of three key components: Information Entropy Clustering, Template Generation, and Chain-of-Thought Merging. Specifically, inspired by information theory Gray (2011), we recognize that different logs encapsulate varying quantities of information, while logs of a similar nature contain comparable amounts. Consequently, we have developed a novel sampling method based on information entropy principles that efficiently clusters characteristic logs by dividing large-scale data into multiple clusters and utilizing efficient sampling and clustering algorithms within those clusters, thereby ensuring that LEMUR remains robust and high-performance in large-scale log scenarios. Then in template generation, we determine the variables and the template in the log based on the information entropy of the token at the same location. In the end, motivated by Chain-of-Thought Wei et al. (2022c), we design a three-hop Chain-of-Thought (infer structure, infer semantics, and infer solution) for merging templates.
@@
We have conducted a comprehensive evaluation on public large-scale log datasets of LogHub He et al. (2020) with seven state-of-the-art log parsing approaches. The results show that Lemur achieves the highest performance than other baselines for the F1 score of grouping and template accuracy. Generally, the main contributions of this work are listed as follows:
@@ 2.1 Information Entropy Clustering
To identify potential log templates from buckets of logs with identical token lengths, we propose a clustering method inspired by information theory Gray (2011), specifically employing Shannon entropy. Unlike conventional clustering which uses random seeds and iterative refinement, our approach selects logs based on their information content, which is ideal for log analysis due to the variability in log messages.
@@
Figure 3: Three-S-hop Chain-of-Thought Merging Method
@@ 2.3 Three-S-hop Chain-of-Thought Merging
HDFS Hadoop Spark Zookeeper BGL HPC Thunderbird Windows Linux Android HealthApp Apache Proxifier OpenSSH OpenStack Mac
Template 14 114 36 50 120 46 149 50 118 166 75 6 8 27 43 341
Average Length 7.44 8.19 8.76 6.30 6.32 3.48 8.51 7.93 8.30 5.40 2.80 6.28 9.35 8.56 9.01 9.17
Various Lengths Proportion 0.25% 0.00% 0.00% 0.55% 0.00% 0.40% 2.25% 0.00% 0.00% 3.85% 0.00% 0.00% 47.35% 0.00% 0.00% 0.35%
Messages 2000 2000 2000 2000 2000 2000 2000 2000 2000 2000 2000 2000 2000 2000 2000 2000
@@
HDFS Hadoop Spark Zookeeper BGL HPC Thunderbird Windows Linux Android HealthApp Apache Proxifier OpenSSH OpenStack Mac
$D$ : =, :, (, ), _ : =, :, , =, ., ., (, ) =, -, : :, = =, :, [, ] =, : (, ) =, : =
$k$ 2 8 6 8 9 9 11 8 25 9 12 12 12 4 20 12
$J_{T}$ 0.7 0.7 0.6 0.9 0.6 0.6 0.4 0.6 0.33 0.7 0.7 0.7 0.7 0.5 0.7 0.7
$\theta$ 2.0 1.7 2.1 2.2 5.5 1.2 4.1 1.1 0.09 3.5 0 0 0.1 0.2 2.3 4.7
Table 2: Transposed hyperparameters setting of Loghub datasets. $D$ is the tokens for word-splitting, $k$ denotes the number of the $S$ in each bucket, $J_{T}$ is the Jaccard similarity threshold for within-bucket merging, and $\theta$ is the entropy threshold for token identification.
@@ 3.1 Implement Details
Experiments are conducted on the most widely-used benchmark datasets published in LogHub (He et al., 2020). More details are available in Table 1.
@@ 3.2 Baselines and Metrics
As for baselines, we choose Drain (He et al., 2017), Spell (Du & Li, 2016), IPLOM (Makanju et al., 2009), ULP (Sedki et al., 2022), Brain (Yu et al., 2023), LogPPT (Le & Zhang, 2023) and LLMParser (Jiang et al., 2023) as our baselines.
@@ 3.3 Main Results
(a) Boxplot of group accuracy on 16 benchmark datasets.
(b) Boxplot of F1 score of group accuracy on 16 benchmark datasets.
Figure 4: Comparison of group accuracy and F1 score on 16 benchmark datasets.
@@
Method Metric HDFS Hadoop Spark Zookeeper BGL HPC Thunderbird Windows Linux Android HealthApp Apache Proxifier OpenSSH OpenStack Mac Average
Lemur w/o CoT FGA 0.999 0.999 1 0.999 0.999 0.999 0.999 1 0.999 0.992 1 1 1 1 1 0.982 0.998
Lemur w/o CoT GA 0.998 0.977 1 0.995 0.989 0.996 0.959 1 0.986 0.956 1 1 1 1 1 0.956 0.988
Lemur FGA 1 1 1 1 1 1 0.999 1 0.999 0.999 1 1 1 1 1 0.984 0.999
Lemur GA 1 1 1 1 1 1 0.982 1 0.988 0.995 1 1 1 1 1 0.977 0.996
Table 5: FGA and GA on LogHub between Lemur with and without CoT.
@@
KMeans Clustering (Sinaga & Yang, 2020) 0.685 0.698
DBSCAN Clustering (Khan et al., 2014) 0.731 0.744
@@
In Figure 4, Lemur shows the robust performance in GA and FGA. The consistency and robustness of Lemur are evident in its high FGA and GA across diverse datasets such as HDFS, Hadoop, Spark, etc. This consistency emphasizes its adaptability and robustness to various log parsing requirements. In conclusion, as an unsupervised LLM-related model, Lemur significantly stands out in the log parsing domain.
@@ 4.2 Additional Ablation Studies on Lemur
To further validate the robustness of Lemur, we conducted ablation studies focusing on the impact of different clustering algorithms and the effectiveness of the bucket grouping strategy. Specifically, we experimented with replacing our clustering mechanism with KMeans (Sinaga & Yang, 2020) and DBSCAN (Khan et al., 2014) and evaluated the performance when disabling the bucket grouping based on the length assumption. These modifications aim to probe the sensitivity of Lemur to these components.
@@ 5.1 Log Parser
In the evolving field of automatic log analysis, crucial for distributed systems and cloud computing, significant progress has been made in log parsing techniques, categorized into frequent pattern mining, clustering, and heuristics rules. Frequent pattern mining is exemplified by SLCT (Vaarandi, 2003) which groups logs based on token frequency, and LogCluster (Vaarandi & Pihelgas, 2015) which removes positional constraints in log grouping. Clustering approaches include LogMine (Hamooni et al., 2016) with its multi-layered clustering system, LKE(Fu et al., 2009) using edit distance and position weighing, SHISO (Mizutani, 2013) improving efficiency through hierarchical clustering, LenMa (Shima, 2016) employing token length vectors, and LPV (Xiao et al., 2020) which uses semantic vectors from word2vec. In the heuristics rules category, IPLOM (Makanju et al., 2009) partitions logs by length and token position, Spell (Du & Li, 2016) approaches parsing as the longest common sequential problem, Drain (He et al., 2017) groups logs by length and prefixes for template updates, and Prefix-Graph (Chu et al., 2021) merges prefix trees into graphs for template generation. Recent advancements have introduced deep learning-based algorithms like Nulog (Nedelkoski et al., 2021) Uniparser (Liu et al., 2022), and LogAP (Rand & Miranskyy, 2021) utilizing comparative learning and machine translation for parsing. However, these deep learning methods face challenges in efficiency and high operational costs due to GPU requirements.
@@ 5.2 Large Language Model
Language modeling using self-supervised learning and large-scale data, significantly enhances various natural language processing tasks. Specifically, pre-training a Transformer decoder (OpenAI, 2023; Vaswani et al., 2017; Yang et al., 2024a; 2025; b; Ouyang et al., 2022a; Wei et al., 2022a; Yang et al., 2023; Nguyen, 2023; Liu et al., 2024; Chai et al., 2024; Zhang et al., 2024b; a; 2025) aids in unconditional text generation. Performance improvements (BertSum, 2023; Yang et al., 2020b; 2022a; 2022b; 2021b; 2021a; a) in diverse tasks have been linked to the enlargement of Pre-training Language Models (PLMs) by increasing model or data size. This has led to the creation of increasingly larger PLMs, such as GPT-3 with 175 billion parameters and PaLM with 540 billion (Anil et al., 2023), guided by the scaling laws of large language models (Aghajanyan et al., 2023). Despite their similar architectures and pre-training tasks, larger PLMs, such as GPT-4 (OpenAI, 2023), exhibit unique behaviors and emergent abilities, excelling in complex tasks. A prime example is ChatGPT, adapting GPT-series LLMs for engaging dialogues, and showcasing advanced conversational skills. Fine-tuning LLMs on various datasets (Wei et al., 2022b) yields promising results, using human or LLM-created prompts for instruction tuning and refining generations. Chain-of-thought prompting (Wei et al., 2022b), where models explain their reasoning for complex problems, and RLHF (Ouyang et al., 2022b), a strategy significantly enhance their performance.